OpenStudy (anonymous):

Please transform the polar equation to an equation in rectangular coordinates. rcsc(theta)=3

OpenStudy (anonymous):

\[(x-3/2)^{2}+y ^{2}=(3/2)^{2}\] correct?

OpenStudy (anonymous):

Anyone? : )

OpenStudy (phi):

r csc(a) = 3 r/sin(a) =3 use x= r*cos(a) y= r*sin(a) x^2+y^2= r^2 so how about this: \[ \frac{r}{r}\cdot \frac{r}{sin(a)}= 3\] \[ \frac{x^2+y^2}{y}=3 \]

OpenStudy (phi):

we can simplify by completing the square (for y), to find the equation of a circle.

OpenStudy (anonymous):

@mikejano, sorry... got called out... yes. your answer is good...

OpenStudy (phi):

no, the center of the circle is at (0,3/2) ( not (3/2,0) )

OpenStudy (anonymous):

yes, phi is correct.